# Solving stochastic climate-economy models: A deep least-squares Monte Carlo approach

### v1

@@
Aleksandar Arandjelović Institute for Statistics and Mathematics, Vienna University of Economics and Business, Austria Pavel V. Shevchenko Date: August 19, 2024
@@
Stochastic versions of recursive integrated climate-economy assessment models are essential for studying and quantifying policy decisions under uncertainty. However, as the number of stochastic shocks increases, solving these models as dynamic programming problems using deterministic grid methods becomes computationally infeasible, and simulation-based methods are needed. The least-squares Monte Carlo (LSMC) method has become popular for solving optimal stochastic control problems in quantitative finance. In this paper, we extend the application of the LSMC method to stochastic climate-economy models. We exemplify this approach using a stochastic version of the DICE model with all five main uncertainties discussed in the literature. To address the complexity and high dimensionality of these models, we incorporate deep neural network approximations in place of standard regression techniques within the LSMC framework. Our results demonstrate that the deep LSMC method can be used to efficiently derive optimal policies for climate-economy models in the presence of uncertainty.
@@ 1 Introduction
The analysis of climate-economy policies is typically performed using Integrated Assessment Models (IAMs) that describe the complex interplay between the climate and the economy via deterministic equations. In order to account for stochastic shocks when finding optimal mitigation policies adapted to climate and economic variables that are evolving stochastically over time, a recursive dynamic programming implementation of integrated assessment models is required. This is a significantly harder computational problem to solve compared to the deterministic case. Seminal contributions to solving IAMs as optimal decision making problems in the presence of uncertainty include Kelly and Kolstad (1999), Kelly and Kolstad (2001), Leach (2007), Traeger (2014), and Cai and Lontzek (2019). All these studies are based on variants of the so-called dynamic integrated climate-economy (DICE) model extended to include stochastic shocks to the economy and climate. The DICE model is one of the three main IAMs (the other two being FUND and PAGE) used by the United States government to determine the social cost of carbon; see Interagency Working Group on Social Cost of Greenhouse Gases (2016). It has been regularly revised over the last three decades, with the first version dating back to Nordhaus et al. (1992). It balances parsimony with realism and is well documented with all published model equations; in addition, its code is publicly available, which is an exception rather than the rule for IAMs. At the same time, it is important to note that IAMs, and the DICE model in particular, have significant limitations (in the model structure and model parameters), which have been criticized and debated in the literature (see the discussions in Ackerman et al. (2009); Pindyck (2017); Grubb et al. (2021); Weitzman (2011)). Despite the criticism, the DICE model has become the iconic typical reference point for climate-economy modeling, and is used in our study.
@@
The original deterministic DICE model is solved as a global optimization problem using the General Algebraic Modeling Language (GAMS) [^1] https://www.gams.com/, a high-level programming language for mathematical modeling. Its stochastic extensions mentioned in the above-mentioned studies require implementations of recursive dynamic programming to find optimal climate policies under uncertainty [^2] If required, the deterministic DICE model can be solved as a recursive dynamic programming problem, too.. This is subject to the curse of dimensionality, and these studies are limited to only one or two stochastic variables. Even in this case, computations take several million core hours on a modern supercomputer (see, for instance, Cai and Lontzek (2019)). Therefore, simulation methods are needed to handle models with many state variables and multiple shocks to reduce the computational burden.
@@
The least-squares Monte Carlo (LSMC) method for solving multi-dimensional stochastic control problems has gained popularity in recent years due to its effectiveness in dealing with high dimensional problems and because it imposes fewer restrictions on the constraints and allows for flexibility in the dynamics of the underlying stochastic processes. The idea is based on simulating random paths of the underlying stochastic variables over time and replacing the conditional expectation of the value function in the Bellman backward recursive solution of the stochastic control problem with an empirical least-squares regression estimate. The transition density of the underlying process is not even required to be known in closed form; one just needs to be able to simulate the underlying processes. The LSMC method was originally developed in Longstaff and Schwartz (2001) and Tsitsiklis and Van Roy (2001). The convergence properties of this method are examined in Belomestny et al. (2010); Belomestny (2011), and Aïd et al. (2014). The LSMC method was originally developed for pricing American options where the state variables are not affected by the control. Later, an extension of the LSMC method with control randomisation was developed in Kharroubi et al. (2014) to handle endogenous state variables (i.e. state variables that are affected by controls). When applied to stochastic control problems that aim to optimize an expected utility, some further extensions are needed as proposed in Andréasson and Shevchenko (2022) and Andréasson and Shevchenko (2024) to achieve a stable and accurate solution.
@@
In this paper, we demonstrate how the LSMC method can be adapted to solve the recursive dynamic programming problem of stochastic IAMs. We exemplify this approach with an application to the DICE model with uncertainties in: (1) the equilibrium temperature sensitivity, (2) the damage function coefficient, (3) the growth rate of total factor productivity, (4) the growth rate of decarbonization, and (5) the equilibrium carbon concentration in the upper strata. These five uncertainties were identified in Nordhaus (2018) as being major sources of uncertainty for the evolution of climate-economic state variables. Typically, polynomial regression is used in LSMC to approximate the corresponding conditional expectations with respect to state variables and controls. However, for models such as the stochastic DICE model, this leads to the need of too many covariates and simulations, making the method not practical. To overcome this problem, we use deep neural network approximations for the required regressions and provide detailed explanations.
@@
The DICE model is a deterministic approach that combines a Ramsey–Cass–Koopmans neoclassical model of economic growth (also known as the Ramsey growth model) with a simple climate model. It involves six state variables (economic capital; temperature in atmosphere and lower oceans; carbon concentration in atmosphere, upper and lower oceans) evolving deterministically in time, two control variables (savings and carbon emission reduction rates) to be determined for each time period of the model, and several exogenous processes (e.g. population size and technology level). The uncertainty about the future of the climate and economy is then typically assessed by treating some model parameters as random variables (because we do not know the exact true value of the key parameters) using a Monte Carlo analysis (see Nordhaus (2018); Gillingham et al. (2015)).
@@
Modeling uncertainty owing to the stochastic nature of the state variables (i.e. owing to the process uncertainty that is present even if we know the model parameters exactly) requires the development and solution of the DICE model as a dynamic model of decision-making under uncertainty, where we calculate the optimal policy response under the assumption of continuing uncertainty throughout the time frame of the model. Few attempts have been made to extend the DICE model to incorporate stochasticity in the underlying state variables and solve it as a recursive dynamic programming problem. For example, Kelly and Kolstad (1999) and Leach (2007) formulated the DICE model with stochasticity in the temporal evolution of temperature, and solved this as a recursive dynamic programming problem. These studies are seminal contributions to the incorporation of uncertainty in the DICE model (although their numerical solution approach is difficult to extend to a higher dimensional space and time-frequency). Cai and Lontzek (2019) formulate DICE as a dynamic programming problem with a stochastic shock on the economy and climate. In addition, Traeger (2014) developed a reduced DICE model with a smaller number of state variables, whereas Lontzek et al. (2015) studied the impact of climate tipping points, and Shevchenko et al. (2022) considered the DICE model with discrete stochastic shocks to the economy. To our best knowledge, the only attempt to solve the stochastic DICE model using an LSMC-type approach is Ikefuji et al. (2020). Their study handles only one uncertainty at a time, and the setup of the regression type Monte Carlo algorithm omits the integration for the conditional expectation in the Bellman equation, assuming the randomness is known in the transition of state variables (in principle, in this case, the required integration can be performed by using deterministic quadrature methods, but this will be subject to the curse of dimensionality).
@@
    We formulate a stochastic version of the DICE model using the sources of uncertainty as identified by Nordhaus (2018). Notably, it does not rely on discretizing the underlying probability distributions that is usually performed in Monte-Carlo type analyses for the sake of model tractability.
@@
    We perform comprehensive numerical experiments and discuss numerical techniques to significantly reduce the computational burden and address several peculiarities of the model. Moreover, we demonstrate how to perform uncertainty quantification (UQ) to understand how uncertainties in the model propagate and affect outputs (such as projections for the evolution of atmospheric temperature).
@@ 2 Model description
In this section, we present the DICE-2016R2 model as a classical example of a recursive climate-economy model. This version of the DICE model was used in Nordhaus (2018). It includes parameter uncertainties in equilibrium temperature sensitivity, the damage function coefficient and the equilibrium carbon concentration in the upper strata, as well as process uncertainties in the growth rate of total factor productivity and the growth rate of decarbonization.
@@
$V=\sup_{\pi}\sum_{t=0}^{\infty}\rho^{t}L_{t}\,u(c_{t}),$ (1)
[…]
The model features six state variables: economic capital $K_{t}$, the concentration of carbon in the atmosphere, the upper oceans, and the lower oceans, $M_{t}=(M_{t}^{\mathrm{AT}},M_{t}^{\mathrm{UP}},M_{t}^{\mathrm{LO}})^{\top}$, and the global mean temperature of the Earth’s surface and the deep oceans, $T_{t}=(T_{t}^{\mathrm{AT}},T_{t}^{\mathrm{LO}})^{\top}$. The evolution of the economic and geophysical sectors is governed by the dynamics described below.
[…]
@@
where $C_{t}$ is total consumption, and $\delta_{K}$ is the rate of depreciation of economic capital.
@@
where $\Phi$ is a coefficient matrix, $E_{t}$ is total $\mathrm{CO}_{2}$ emissions (in billions of tons per year), and $\beta$ is the conversion factor of $\mathrm{CO}_{2}$ mass into the equivalent mass of carbon. Emissions $E_{t}$ are equal to uncontrolled industrial emissions, given by a level of carbon intensity (see Subsection 2.1) $\sigma_{t}$ times gross output, reduced by the emission reduction rate $\mu_{t}$, plus exogenous land-use emissions $\widetilde{E}_{t}$, i.e. $E_{t}=\sigma_{t}(1-\mu_{t})Y_{t}+\widetilde{E}_{t}$.
@@
where $V_{t}$ denotes the value function at time $t$, and $\beta$ represents the $\mathrm{CO}_{2}$ to carbon mass transformation coefficient.
@@
$\Lambda_{t}=550(1-0.025)^{t}/(1000\theta_{2})\sigma_{t}\mu_{t}^{\theta_{2}}$
@@
$\alpha=1.45$, $\beta=1/3.666$, $\gamma=0.3$, $\rho=0.015$, $\delta_{K}=0.1$

### v2

# disposition

@@
Aleksandar Arandjelović Department of Mathematics, ETH Zurich, Switzerland Institute for Statistics and Mathematics, Vienna University of Economics and Business, Austria Pavel V. Shevchenko Date: 25. Dezember 2025
@@
Stochastic versions of recursive integrated climate-economy assessment models are essential for studying and quantifying policy decisions under uncertainty. However, as the number of state variables and stochastic shocks increases, solving these models via deterministic grid-based dynamic programming (e.g., value-function iteration / projection on a discretized grid over continuous state variables, typically coupled with discretized shocks) becomes computationally infeasible, and simulation-based methods are needed. The least-squares Monte Carlo (LSMC) method has become popular for solving optimal stochastic control problems in quantitative finance. In this paper, we extend the application of the LSMC method to stochastic climate-economy models. We exemplify this approach using a stochastic version of the DICE model with five key uncertainty sources highlighted in the literature. To address the complexity and high dimensionality of these models, we incorporate deep neural network approximations in place of standard regression techniques within the LSMC framework. Our results demonstrate that the deep LSMC method can be used to efficiently derive optimal policies for climate-economy models in the presence of uncertainty.
@@ 1 Introduction
The analysis of climate-economy policies is typically performed using Integrated Assessment Models (IAMs) that describe the complex interplay between the climate and the economy via deterministic equations. In order to account for stochastic shocks when finding optimal mitigation policies adapted to climate and economic variables that are evolving stochastically over time, a recursive dynamic programming implementation of integrated assessment models is required. This is a significantly harder computational problem to solve compared to the deterministic case. Seminal contributions to solving IAMs as optimal decision-making problems in the presence of uncertainty include kelly1999, kelly2001solving, leach2007, traeger2014, and cai2019. All these studies are based on variants of the so-called dynamic integrated climate-economy (DICE) model extended to include stochastic shocks to the economy and climate. The DICE model is one of the three main IAMs (the other two being FUND and PAGE) used by the United States government to determine the social cost of carbon; see scc2016technical. It has been regularly revised over the last three decades, with the first version dating back to nordhaus1992dice. It balances parsimony with realism and is well documented with all published model equations; in addition, its code is publicly available, which is an exception rather than the rule for IAMs. At the same time, it is important to note that IAMs, and the DICE model in particular, have significant limitations (in the model structure and model parameters), which have been criticized and debated in the literature (see the discussions in ackerman2009limitations; pindyck2017use; grubb2021modeling; weitzman2011fat). Recent work has also raised a more specific concern about the benchmark calibration of the reduced-form climate emulator used in DICE-2016: dietz2021climatedynamics and folini2025climate document that the standard DICE-2016 calibration can deviate from climate-science benchmarks for key aspects of the transient warming response (and related heat-uptake dynamics). In this paper, we retain the DICE-2016R2 calibration as a widely used benchmark to facilitate comparison with the IAM literature and to keep the focus on the numerical method; our quantitative results should therefore be interpreted conditional on this benchmark climate block. Despite the criticism, the DICE model has become the iconic typical reference point for climate-economy modeling and is used in our study.
@@
The original deterministic DICE model is solved as a global optimization problem using the General Algebraic Modeling Language (GAMS) [^1] https://www.gams.com/, a high-level programming language for mathematical modeling. Its stochastic extensions mentioned in the above-mentioned studies require implementations of recursive dynamic programming to find optimal climate policies under uncertainty [^2] If required, the deterministic DICE model can be solved as a recursive dynamic programming problem too.. This is subject to the curse of dimensionality, and these studies are limited to only one or two stochastic variables. Even in this case, computations take several million core hours on a modern supercomputer (see, for instance, cai2019). Therefore, simulation methods are needed to handle models with many state variables and multiple shocks to reduce the computational burden.
@@
The least-squares Monte Carlo (LSMC) method for solving multi-dimensional stochastic control problems has gained popularity in recent years due to its effectiveness in dealing with high dimensional problems and because it imposes fewer restrictions on the constraints and allows for flexibility in the dynamics of the underlying stochastic processes. The idea is based on simulating random paths of the underlying stochastic variables over time and replacing the conditional expectation of the value function in the Bellman backward recursive solution of the stochastic control problem with an empirical least-squares regression estimate. The transition density of the underlying process is not even required to be known in closed form; one just needs to be able to simulate the underlying processes. The LSMC method was originally developed in longstaff2001valuing and tsitsiklis2001regression. The convergence properties of this method are examined in belomestny2010regression; belomestny2011pricing, and aid2014probabilistic. The LSMC method was originally developed for pricing American options where the state variables are not affected by the control. Later, an extension of the LSMC method with control randomisation was developed in kharroubi2014numerical to handle endogenous state variables (i.e. state variables that are affected by controls). When applied to stochastic control problems that aim to optimize an expected utility, some further extensions are needed as proposed in andreasson2022 and andreasson2024optimal to achieve a stable and accurate solution.
Our numerical backbone builds on the bias-corrected endogenous-state least-squares Monte Carlo recursion for multi-period expected-utility control developed by andreasson2022. The present paper does not claim a new Bellman recursion; rather, it contributes the methodological and implementation extensions required to make this framework practically usable for stochastic integrated assessment models. In particular, relative to andreasson2022 we: (i) couple the bias-corrected recursion with deep neural-network regressors for continuation-value conditional expectations in a moderately high-dimensional climate-economy state; (ii) introduce numerically critical stabilization choices in the stochastic DICE setting (post-decision states, normalization of rapidly growing economic variables, and low-discrepancy sampling); and (iii) demonstrate how a single backward recursion can be reused for uncertainty quantification (e.g. distributions of the social cost of carbon and variance-based Sobol’ indices) via forward simulation under the resulting optimal feedback policy.
Regression approximations of conditional expectations have a long tradition in computational macroeconomics. In particular, the parameterized expectations approach of den1990solving computes expectations appearing in equilibrium conditions (typically Euler equations) via forward in time simulation with regression; duffy2001approximating extends this idea using neural networks, and valaitis2024machine revisits related projection ideas from a modern machine-learning perspective in macro-finance models. Our method is different: backward in time, we approximate the continuation value conditional expectation in the Bellman recursion using least-squares Monte Carlo with control randomization and deep neural network regressors, and then recover the optimal control by a separate numerical maximization step. This makes the method directly applicable to controlled, high-dimensional climate-economy dynamics with multiple shocks and policy constraints, while keeping the learning step (conditional expectation) and the optimization step (policy choice) cleanly separated.
@@
In this paper, we demonstrate how the LSMC method can be adapted to solve the recursive dynamic programming problem of stochastic IAMs. Our goal is not to deliver a new DICE specification or to claim state-of-the-art climate-economic calibration; rather, we use the standard DICE-2016R2 setup as a well-documented test-case to demonstrate how the proposed deep LSMC methodology scales to multiple sources of uncertainty. We exemplify this approach with an application to the DICE model with uncertainties in: (1) equilibrium climate sensitivity, (2) the damage function coefficient, (3) the growth rate of total factor productivity, (4) the growth rate of decarbonization, and (5) the equilibrium carbon concentration in the upper strata, where (3)–(4) enter as stochastic process shocks and (1)–(2), (5) are treated as time-invariant (initial) parameter uncertainty. These five uncertainties were identified in nordhaus2018 as being major sources of uncertainty for the evolution of climate-economic state variables. Typically, polynomial regression is used in LSMC to approximate the corresponding conditional expectations with respect to state variables and controls. However, for models such as the stochastic DICE model, this leads to the need for too many covariates and simulations, making the method not practical. To overcome this problem, we use deep neural network approximations for the required regressions and provide detailed explanations. In addition to computing optimal feedback policies, we use the deep LSMC solution to perform uncertainty quantification for key quantities of interest (e.g., the social cost of carbon and temperature trajectories) and variance-based global sensitivity analysis via Sobol’ indices; see Section 4.
@@
The DICE model is a deterministic approach that combines a Ramsey–Cass–Koopmans neoclassical model of economic growth (also known as the Ramsey growth model) with a simple climate model. It involves six state variables (economic capital; temperature in atmosphere and lower oceans; carbon concentration in atmosphere, upper and lower oceans) evolving deterministically in time, two control variables (savings and carbon emission reduction rates) to be determined for each time period of the model, and several exogenous processes (e.g. population size and technology level). The uncertainty about the future of the climate and economy is then typically assessed by treating some model parameters as random variables (because we do not know the exact true value of the key parameters) using a Monte Carlo analysis (see nordhaus2018; gillingham2015modeling).
@@
Modeling uncertainty owing to the stochastic nature of the state variables (i.e. owing to the process uncertainty that is present even if we know the model parameters exactly) requires the development and solution of the DICE model as a dynamic model of decision-making under uncertainty, where we calculate the optimal policy response under the assumption of continuing uncertainty throughout the time frame of the model. Few attempts have been made to extend the DICE model to incorporate stochasticity in the underlying state variables and solve it as a recursive dynamic programming problem. For example, kelly1999 and leach2007 formulated the DICE model with stochasticity in the temporal evolution of temperature, and solved this as a recursive dynamic programming problem. These studies are seminal contributions to the incorporation of uncertainty in the DICE model (although their numerical solution approach is difficult to extend to a higher dimensional space and time-frequency). cai2019 formulate DICE as a dynamic programming problem with a stochastic shock on the economy and climate. In addition, traeger2014 developed a reduced DICE model with a smaller number of state variables, whereas lontzek2015stochastic studied the impact of climate tipping points, and shevchenko2022impact considered the DICE model with discrete stochastic shocks to the economy. To the best of our knowledge, the only attempt to solve the stochastic DICE model using an LSMC-type approach is ikefuji2020expected. Their study handles only one uncertainty at a time, and the setup of the regression type Monte Carlo algorithm omits the integration for the conditional expectation in the Bellman equation, assuming the randomness is known in the transition of state variables (in principle, in this case, the required integration can be performed by using deterministic quadrature methods, but this will be subject to the curse of dimensionality).
@@
    We formulate a stochastic version of the DICE model using the sources of uncertainty as identified by nordhaus2018. Notably, it does not rely on discretizing the underlying probability distributions, as is usually done in Monte-Carlo-type analyses for the sake of model tractability.
@@
    We perform comprehensive numerical experiments and discuss uncertainty quantification (UQ) for key quantities of interest. Our variance-based global sensitivity analysis (GSA) using Sobol’ indices is closely related to modern UQ/GSA approaches in economics and computational economics (harenberg2019uqgsa; scheidegger2019mlhd) and, in particular, to the DICE sensitivity analysis of miftakhova2021gsa.
The main contribution of this paper is methodological: we propose a deep least-squares Monte Carlo solver for recursive climate-economy stochastic control problems with endogenous state dynamics, inequality-constrained policies, and multiple uncertainty sources. [^3] The code used for the numerical experiments is available upon request. We use the DICE-2016R2 calibration and the five uncertainty blocks reported in nordhaus2018 as a transparent benchmark to facilitate comparison with the IAM literature, not because we view this particular uncertainty specification as uniquely policy-relevant or definitive. The deep LSMC algorithm only requires a simulatable one-step transition map; therefore, alternative uncertainty blocks, recalibrations, or climate-emulator variants can be incorporated without conceptual changes to the backward-recursion framework. For practical guidance on when deep LSMC is preferable to grid / VFI / projection approaches (and when it is not), see Section 3.4.
While the present paper is primarily methodological, the stochastic-DICE output distributions produced by our framework (e.g. temperature paths, the SCC, and policy trajectories under uncertainty) can be used as inputs to climate-finance and actuarial risk assessments. For instance, arandjelovicShevchenko2025netzero quantify uncertainty in the SCC and related tail-risk measures under net-zero scenarios, and arandjelovicShevchenko2025reserves develop a stress-testing framework for life-insurance reserving driven by stochastic-DICE temperature projections.
@@ 2 Model description
In this section, we present the DICE-2016R2 model as a classical example of a recursive climate-economy model. This version of the DICE model was used in nordhaus2018. It includes parameter uncertainties in equilibrium climate sensitivity, the damage function coefficient and the carbon-cycle coefficient (equilibrium carbon concentration in the biosphere/upper ocean strata), as well as process uncertainties in the growth rate of total factor productivity and the growth rate of decarbonization.
@@
$V_{0}=\sup_{\pi}\sum_{t=0}^{\infty}\rho^{t}L_{t}\,u\left(\frac{C_{t}}{L_{t}}\right),$ (1)
where $\rho\in(0,1)$ is the $\Delta$-step discount factor (in the DICE-2016R2 calibration the annual pure rate of time preference is $0.015$, hence $\rho=(1+0.015)^{-\Delta}$), $L_{t}$ is world population, $C_{t}$ is total consumption, and the time index $t=0,1,\ldots$ corresponds to $\Delta=5$-year steps. In the numerical implementation (as in standard DICE codes), we approximate this by a long but finite horizon with a terminal reward; see the discussion around Table 1 and Section 4. We parameterize $\pi_{t}=(c_{t},\mu_{t})$, where $c_{t}\in[0,1]$ is the consumption share of net output $Q_{t}$ (so that $C_{t}=c_{t}Q_{t}$ and per-capita consumption equals $C_{t}/L_{t}$) and $\mu_{t}\in[0,1]$ is the industrial-emissions mitigation rate. The utility function $u$ has constant relative risk aversion, $u(c)=\frac{c^{1-\alpha}-1}{1-\alpha}$, with risk-aversion parameter $\alpha\geq 0$ (the case $\alpha=1$ corresponds to logarithmic utility).
Remark 2.1.
The baseline DICE specification uses CRRA utility. While convenient and standard in the DICE benchmark, CRRA ties together the coefficient of relative risk aversion and the elasticity of intertemporal substitution (IES), which can attenuate precautionary motives and thereby mute the observable impact of uncertainty on optimal policies and the social cost of carbon (SCC). To separate risk aversion and intertemporal substitution, one may instead adopt recursive preferences such as Epstein–Zin utility (epstein1989substitution; weil1990nonexpected), as used in stochastic IAM analyses including cai2019. Our deep LSMC approach extends directly to this case because the Bellman recursion remains backward and regression-based: the continuation step requires approximating conditional expectations of a (power-transformed) continuation value.
The deterministic DICE core features six endogenous state variables: economic capital $K_{t}$, the concentration of carbon in the atmosphere, the upper oceans, and the lower oceans, $M_{t}=(M_{t}^{\mathrm{AT}},M_{t}^{\mathrm{UP}},M_{t}^{\mathrm{LO}})^{\top}$, and the global mean temperature of the Earth’s surface and the deep oceans, $T_{t}=(T_{t}^{\mathrm{AT}},T_{t}^{\mathrm{LO}})^{\top}$. In our stochastic specification (Subsection 2.1), the Markov state used in the numerical recursion is augmented by the realized exogenous levels $(A_{t},\sigma_{t})$ and by the three uncertain structural parameters (drawn at $t=0$ and held fixed along a trajectory); see Section 4 for the resulting 11-dimensional state vector used in the numerical study. The evolution of the economic and geophysical sectors is governed by the dynamics described below.
The economic system: Gross output is modeled by a Cobb–Douglas production function of capital, labor, and technology, $Y_{t}=A_{t}K_{t}^{\gamma}L_{t}^{1-\gamma}$, where $\gamma\in(0,1)$ and $1-\gamma$ are the output elasticities of capital and labor, respectively. Here, $A_{t}$ denotes total factor productivity (see Subsection 2.1), representing technological progress and efficiency improvements over time.
@@
where $\delta_{K}$ is the rate of depreciation of economic capital. In terms of the consumption-share control $c_{t}$ defined above, total consumption is $C_{t}=c_{t}Q_{t}$, so that investment equals $(1-c_{t})Q_{t}$ and the capital update can be written as $K_{t+1}=(1-\delta_{K})^{\Delta}K_{t}+\Delta\times(1-c_{t})Q_{t}$.
@@
where $\Phi$ is a coefficient matrix, $E_{t}$ is total $\mathrm{CO}_{2}$ emissions (in billions of tons per year), and $\beta$ is conversion factor of $\mathrm{CO}_{2}$ mass into the equivalent mass of carbon. Emissions $E_{t}$ are equal to uncontrolled industrial emissions, given by a level of carbon intensity (see Subsection 2.1) $\sigma_{t}$ times gross output, reduced by the emission reduction rate $\mu_{t}$, plus exogenous land-use emissions $\widetilde{E}_{t}$, i.e. $E_{t}=\sigma_{t}(1-\mu_{t})Y_{t}+\widetilde{E}_{t}$.
@@
Remark 2.2.
Recent work by dietz2021climatedynamics and folini2025climate shows that the benchmark climate block in DICE-2016/DICE-2016R2 does not fully reproduce the behavior of more detailed Earth-system models for key physical benchmarks, including aspects of the transient temperature response and related ocean heat-uptake dynamics. It was demonstrated in folini2025climate that the functional form of the climate variable equations in DICE-2016 is fit for purpose and can be re-calibrated to match the benchmark data. In this paper we retain the DICE-2016R2 calibration as a widely used benchmark in order to facilitate comparison with the existing IAM literature and to focus on the numerical contribution. Accordingly, quantitative results should be interpreted conditional on this benchmark climate block. At the same time, the deep LSMC method developed below is not tied to this particular emulator: it requires only a simulatable one-step transition map and can therefore be combined with alternative (climate-science-consistent) climate emulators, including re-calibrated DICE variants as in folini2025climate.
@@
Remark 2.3.
For readers more familiar with continuous-time stochastic control, the discrete-time DICE transition equations can be viewed as arising from sampling (and discretising) an underlying continuous-time controlled system on calendar time $s$ and then evaluating it at a grid with step size $\Delta=5$ years.
For instance, a continuous-time capital-accumulation dynamics of the form
$\mathrm{d}K_{s}=\big(Q_{s}-C_{s}-\delta_{K}K_{s}\big)\,\mathrm{d}s\;+\;\nu_{K}(X_{s},\pi_{s})\,\mathrm{d}W_{s},$
(with state vector $X_{s}$, Brownian motion $W_{s}$, and $\nu_{K}\equiv 0$ in the deterministic DICE-2016R2 calibration) yields, after holding controls approximately constant over each $\Delta$-year period and discretising in time, a one-step update of the form used in the discrete-time model. The carbon-cycle and temperature blocks admit analogous continuous-time representations, with discrete-time innovations interpretable as aggregated shocks over a $\Delta$-year period.
We retain the standard $\Delta$-step DICE-2016R2 specification in order to preserve the benchmark calibration and facilitate comparison with the IAM literature. The deep LSMC method developed in Section 3 only requires that the resulting one-step transition map can be simulated, and therefore applies equally to time-discretised continuous-time climate-economy models; see also cai2012 for a continuous-time integrated-assessment formulation.
@@
where $V_{t}$ denotes the value function at time $t$, and $\beta$ represents the $\mathrm{CO}_{2}$ to carbon mass transformation coefficient. Here $K_{t}$ is measured in trillions of 2010 USD and $M_{t}^{\mathrm{AT}}$ in gigatons of carbon, so $(\partial V_{t}/\partial M_{t}^{\mathrm{AT}})/(\partial V_{t}/\partial K_{t})$ has units “trillion USD per GtC”; multiplying by $1000$ converts to USD per tC, and $\beta=1/3.666$ converts from tC to tCO2.
@@
$\Lambda_{t}=550(1-0.025)^{t}\sigma_{t}\mu_{t}^{\theta_{2}}/(1000\theta_{2})$
@@
$\alpha=1.45$, $\beta=1/3.666$, $\gamma=0.3$, $\rho=(1+0.015)^{-\Delta}$, $\delta_{K}=0.1$